# Characterisation of distal actions of automorphisms on the space of one-parameter subgroups of Lie groups

### v2

@@ 4. Proofs of the main results
Now we show that $(1)\implies(7)$. Suppose every $T\in{\mathcal{H}}$ acts distally on ${\rm Sub}^{p}_{G}$. Let $T\in{\mathcal{H}}$ be fixed. By Theorem 1.1 (1), $T\in{\mathcal{B}}{\mathcal{D}}$, for some compact group ${\mathcal{B}}\subset{\rm GL}(d,{\mathbb{R}})$. Let ${\mathcal{H}}_{1}:=\overline{[{\mathcal{H}},{\mathcal{H}}]}$. Then ${\mathcal{H}}_{1}\subset{\mathcal{H}}$ as the latter is closed. Now let $T\in{\mathcal{H}}_{1}$. Then $\det T=1$. From above we have that $T=\phi\,\circ\,r\,{\rm Id}$ for some $\phi\in{\mathcal{B}}$ and some $r\in{\mathbb{R}}\setminus\{0\}$. Since ${\mathcal{B}}$ is a compact group, $\det\phi=\pm 1$, and hence $r=\pm 1$. Therefore, $T$ generates a relatively compact group, for every $T\in{\mathcal{H}}_{1}$. As ${\mathcal{H}}_{1}$ is a closed subgroup of ${\rm GL}(d,{\mathbb{R}})$ by Theorem 1.1 of [17] or Proposition 2 of [39], ${\mathcal{H}}_{1}$ is compact. Now ${\mathcal{H}}_{1}$ is a compact normal subgroup of ${\mathcal{H}}$. Let $N({\mathcal{H}}_{1})$ be the normaliser of ${\mathcal{H}}_{1}$ in ${\rm GL}(d,{\mathbb{R}})$. Then $N({\mathcal{H}}_{1})$ is a closed algebraic subgroup of ${\rm GL}(d,{\mathbb{R}})$ and ${\mathcal{H}},{\mathcal{D}}\subset N({\mathcal{H}}_{1})$. Moreover, $N({\mathcal{H}}_{1})\subset N({\mathcal{H}}_{1}{\mathcal{D}})$. Being algebraic, $N({\mathcal{H}}_{1})$ has finitely many connected components, and hence so does $N({\mathcal{H}}_{1})/{\mathcal{H}}_{1}{\mathcal{D}}$. In particular any closed abelian subgroup of $N({\mathcal{H}}_{1})/{\mathcal{H}}_{1}{\mathcal{D}}$ is compactly generated (cf. [24]). Let $\rho:N({\mathcal{H}}_{1})\to N({\mathcal{H}}_{1})/{\mathcal{H}}_{1}{\mathcal{D}}$ be the natural projection. Then $\rho({\mathcal{H}})$ is abelian, and hence so is $\overline{\rho({\mathcal{H}})}$. Therefore, $\overline{\rho({\mathcal{H}})}$ is compactly generated, and it has a unique maximal compact subgroup, say ${\mathcal{C}}$.
@@
Suppose (1) holds; i.e. every inner automorphism of $G$ acts distally on ${\rm Sub}^{p}_{G}$. If $G$ is either compact or a vector group, then (5) holds trivially. Now suppose $G$ neither compact nor a vector group. Since every inner automorphism acts trivially on the center of $G$, we have by Theorem 1.2 that it acts distally on $G$. Therefore, $G$ is distal (cf. [36]), i.e. the action of ${\rm Inn}(G)$ on $G$ is distal. It follows from Theorem 9 of [36] and Corollary 2.1 (ii) of [27] that $G/R$ is compact, where $R$ is the radical of $G$. Let $N$ be the nilradical of $G$ and let $x\in N$. Then ${\rm inn}(x)$ is unipotent, and by Theorem 1.3, ${\rm inn}(x)={\rm Id}$. This implies that $x$ is central in $G$ for every $x\in N$. Then $N$ is abelian and central in $G$. Now for the radical $R$, we know that $[R,R]\subset N$ is central in $G$, and hence $R$ is nilpotent and it is same as $N$ and it is abelian and central in $G$. Now $G/N$ is compact, where $N$ is central in $G$, i.e. $N={\mathbb{R}}^{n}\times M$, for some $n\in{\mathbb{N}}$, where $M$ is compact. Here, $V:={\mathbb{R}}^{n}$ is a vector subgroup which is central. By Lemma 3.7 of [25], $G=K\times{\mathbb{R}}^{n}$, where $K$ is the maximal compact subgroup of $G$. Thus (5) holds. ∎
The distal actions of automorphisms $T$ of a connected Lie group $G$ on ${\rm Sub}^{a}_{G}$ were studied earlier if $G$ is in class ${\mathcal{C}^{\prime}}$ or $T$ is unipotent. Here we have studied a more general case when $T$ acts distally on the maximal central torus of $G$ and the action is considered on ${\rm Sub}^{p}_{G}$, a smaller subspace of ${\rm Sub}^{a}_{G}$. More generally, we have also studied the action of subgroups ${\mathcal{H}}$ of ${\rm Aut}(G)$ on ${\rm Sub}^{p}_{G}$ under a condition that the ${\mathcal{H}}$-action on the maximal central torus of $G$ is distal. This latter condition is satisfied if ${\mathcal{H}}$ is contained in ${\rm Aut}(G)^{0}$ or more generally, if it is contained in an almost algebraic subgroup of ${\rm Aut}(G)$.
@@
Let $\mathbb{H}$ be the 3-dimensional Heisenberg group of $3\times 3$ strictly upper triangular matrices. There is a canonical action of $SL(2,{\mathbb{R}})$ on $\mathbb{H}$ which fixes elements of its center, which is isomorphic to ${\mathbb{R}}$. Let $N=\mathbb{H}/D$, where $D$ is a discrete central subgroup isomorphic to ${\mathbb{Z}}$. Then $N$ is a 2-step nilpotent group and the ${\rm SL}(2,{\mathbb{R}})$-action on $\mathbb{H}$ carries over to the action on $N$. Let $G=SO(2,{\mathbb{R}})\ltimes N$, the Walnut group. Since the center of $G$ is isomorphic to ${\mathbb{S}}^{1}$, ${\rm Aut}(G)$ acts distally on it. Thus Theorems 1.2 and 1.4 hold for the action of any $T\in{\rm Aut}(G)$ and Theorem 1.5 holds for any subgroup $H$ of ${\rm Aut}(G)$.

### v3

@@ 4. Proofs of the main results
Now we show that $(1)\implies(7)$. Suppose every $T\in{\mathcal{H}}$ acts distally on ${\rm Sub}^{p}_{G}$. Let $T\in{\mathcal{H}}$ be fixed. By Theorem 1.1 (1), $T\in{\mathcal{B}}{\mathcal{D}}$, for some compact group ${\mathcal{B}}\subset{\rm GL}(d,{\mathbb{R}})$. Let ${\mathcal{H}}_{1}:=\overline{[{\mathcal{H}},{\mathcal{H}}]}$. Then ${\mathcal{H}}_{1}\subset{\mathcal{H}}$ as the latter is closed. Now let $T\in{\mathcal{H}}_{1}$. Then $\det T=1$. From above we have that $T=\phi\,\circ\,r\,{\rm Id}$ for some $\phi\in{\mathcal{B}}$ and some $r\in{\mathbb{R}}\setminus\{0\}$. Since ${\mathcal{B}}$ is a compact group, $\det\phi=\pm 1$, and hence $r=\pm 1$. Therefore, $T$ generates a relatively compact group, for every $T\in{\mathcal{H}}_{1}$. As ${\mathcal{H}}_{1}$ is a closed subgroup of ${\rm GL}(d,{\mathbb{R}})$ by Theorem 1.1 of [17] or Proposition 2 of [39], ${\mathcal{H}}_{1}$ is compact. Let $N({\mathcal{H}}_{1})$ be the normaliser of ${\mathcal{H}}_{1}$ in ${\rm GL}(d,{\mathbb{R}})$. Then $N({\mathcal{H}}_{1})$ is a closed algebraic subgroup of ${\rm GL}(d,{\mathbb{R}})$ and ${\mathcal{H}},{\mathcal{D}}\subset N({\mathcal{H}}_{1})$. Moreover, $N({\mathcal{H}}_{1})\subset N({\mathcal{H}}_{1}{\mathcal{D}})$. Being algebraic, $N({\mathcal{H}}_{1})$ has finitely many connected components, and hence so does $N({\mathcal{H}}_{1})/{\mathcal{H}}_{1}{\mathcal{D}}$. In particular any closed abelian subgroup of $N({\mathcal{H}}_{1})/{\mathcal{H}}_{1}{\mathcal{D}}$ is compactly generated (cf. [24]). Let $\rho:N({\mathcal{H}}_{1})\to N({\mathcal{H}}_{1})/{\mathcal{H}}_{1}{\mathcal{D}}$ be the natural projection. Then $\rho({\mathcal{H}})$ is abelian, and hence so is $\overline{\rho({\mathcal{H}})}$. Therefore, $\overline{\rho({\mathcal{H}})}$ is compactly generated, and it has a unique maximal compact subgroup, say $\mathcal{L}$.
@@
Suppose (1) holds; i.e. every inner automorphism of $G$ acts distally on ${\rm Sub}^{p}_{G}$. If $G$ is either compact or a vector group, then (5) holds trivially. Now suppose $G$ neither compact nor a vector group. Since every inner automorphism acts trivially on the center of $G$, we have by Theorem 1.2 that it acts distally on $G$. Therefore, $G$ is distal (cf. [36]), i.e. the action of ${\rm Inn}(G)$ on $G$ is distal. It follows from Theorem 9 of [36] and Corollary 2.1 (ii) of [27] that $G/R$ is compact, where $R$ is the radical of $G$. Let $N$ be the nilradical of $G$ and let $x\in N$. Then ${\rm inn}(x)$ is unipotent, and by Theorem 1.3, ${\rm inn}(x)={\rm Id}$. This implies that $x$ belongs to the center of $G$ for every $x\in N$. Thus $N$ is abelian and central in $G$. Now for the radical $R$, we know that $[R,R]\subset N$ is central in $G$, and hence $R$ is nilpotent and it is same as $N$ and it is abelian and central in $G$. Now $G/N$ is compact, where $N$ is central in $G$, i.e. $N={\mathbb{R}}^{n}\times M$, for some $n\in{\mathbb{N}}$, where $M$ is compact. Here, $V:={\mathbb{R}}^{n}$ is a vector subgroup which is central. By Lemma 3.7 of [25], $G={\mathbb{R}}^{n}\times K$, where $K$ is the maximal compact subgroup of $G$. Thus (5) holds. ∎
The distal actions of automorphisms $T$ of a connected Lie group $G$ on ${\rm Sub}^{a}_{G}$ were studied earlier if $G$ is in class ${\mathcal{C}^{\prime}}$ or $T$ is unipotent. Here we have studied a more general case when $T$ acts distally on the maximal central torus of $G$ and the action is considered on ${\rm Sub}^{p}_{G}$, a smaller subspace of ${\rm Sub}^{a}_{G}$. More generally, we have also studied the actions of subgroups ${\mathcal{H}}$ of ${\rm Aut}(G)$ on ${\rm Sub}^{p}_{G}$ under a condition that the ${\mathcal{H}}$-action on the maximal central torus of $G$ is distal. This latter condition is satisfied if ${\mathcal{H}}$ is contained in ${\rm Aut}(G)^{0}$ or more generally, if it is contained in an almost algebraic subgroup of ${\rm Aut}(G)$.
@@
Let $\mathbb{H}$ be the 3-dimensional Heisenberg group consisting of  $\,3\times 3$ real strictly upper triangular matrices. There is a canonical action of $SL(2,{\mathbb{R}})$ on $\mathbb{H}$ which fixes elements of its center, which is isomorphic to ${\mathbb{R}}$. Let $N=\mathbb{H}/D$, where $D$ is a discrete central subgroup isomorphic to ${\mathbb{Z}}$. Then $N$ is a connected 2-step nilpotent group and the ${\rm SL}(2,{\mathbb{R}})$-action on $\mathbb{H}$ carries over to the action on $N$. Let $G=SO(2,{\mathbb{R}})\ltimes N$, the Walnut group. Since the center of $G$ is isomorphic to ${\mathbb{S}}^{1}$, ${\rm Aut}(G)$ acts distally on it. Thus Theorems 1.2 and 1.4 hold for the action of any $T\in{\rm Aut}(G)$ and Theorem 1.5 holds for any subgroup ${\mathcal{H}}$ of ${\rm Aut}(G)$.